OpenStudy (anonymous):

Find the slope of the line passing through (4, 9) and (-2, 9). 0 undefined Find the slope of the line passing through (4, 9) and (-2, 9). 0 undefined @Mathematics

OpenStudy (anonymous):

slope = rise / run = y2 - y1 / x2 - x1 = 9 - 9 / -2 - 4 = 0